Jenny is training for a race. On Day 1, she ran5 miles, which was 1/3 the distance she ran on Day 3, and 1/2 the distance she ran on Day 2. How many miles did she run over the 3-day period?
Answer by lwsshak3(11628)

Jenny is training for a race. On Day 1, she ran5 miles, which was 1/3 the distance she ran on Day 3, and 1/2 the distance she ran on Day 2. How many miles did she run over the 3-day period?
***
distance Jenny ran over the 3-day period:
1st day=5 miles
2nd day=2*5=10
3rd day=3*5=15
3-day total=30 miles